OpenStudy (anonymous):

what is your x-intercept when x=2 and b=4 what is your x-intercept when x=2 and b=4 @Mathematics

OpenStudy (saifoo.khan):

isnt it already given?

OpenStudy (anonymous):

>D

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I meant m=2

OpenStudy (saifoo.khan):

y = mx + b y = 2x + 4

hero (hero):

y = mx + b y = 2x + 4

OpenStudy (anonymous):

Give me a medal pleaswe

hero (hero):

-2

OpenStudy (anonymous):

Hero!

hero (hero):

x int = -2

OpenStudy (saifoo.khan):

when x intercepts, y = 0 y = 2x + 4 0 = 2x + 4 -4 = 2x x = -4/2 x = -2 x intercept = (-2,0)

OpenStudy (anonymous):

mm thank you :)

OpenStudy (saifoo.khan):

no problem.

hero (hero):

jurneelynn, I provided a better explanation for your previous problem

OpenStudy (saifoo.khan):

your wish.

hero (hero):

Shut up Saifoo

OpenStudy (saifoo.khan):

U shut up. :P

hero (hero):

I told you to shut up first

OpenStudy (saifoo.khan):

that dosnt matter.

OpenStudy (anonymous):

LOLOLOLOLOLOL

OpenStudy (anonymous):